OpenStudy (ray1998):

Please help? Will give fan and medal. Write the arithmetic series in summation notation. 5+10+15+20+25+30+35 =? Thank you in advance!!

rishavraj (rishavraj):

so u got all the values....u just need to plug them in the equation :))

OpenStudy (ray1998):

Okay. So how? :)

rishavraj (rishavraj):

\[S_n = \frac{ n }{ 2 }(a_1 + a_n)\] \[a_1 = 5~~and~~a_n = 35~~~and~~n = no ~~of~~ terms ...\]

rishavraj (rishavraj):

count the no of terms :))

OpenStudy (ray1998):

Okay, thank you!

rishavraj (rishavraj):

u welcome :)

OpenStudy (ray1998):

Is it, n = 140/s?

rishavraj (rishavraj):

n is the no of terms ... so n = 7

rishavraj (rishavraj):

so it means \[S_7 = \frac{ 7 }{ 2 }(5 + 35)\]

rishavraj (rishavraj):

@Ray1998 oops did u mean S_n = 140 ??? i mean the sum??

OpenStudy (ray1998):

Yes. :)

rishavraj (rishavraj):

no no there would be 7

rishavraj (rishavraj):

140 is the value of sum of tht sequence

OpenStudy (ray1998):

7 goes where I have the 140?

OpenStudy (ray1998):

And the 140 on the side?

rishavraj (rishavraj):

tbh not sure about the ...bt 7 goes where u plugged 140

OpenStudy (ray1998):

Okay. Thank you :)

rishavraj (rishavraj):

u welcome :))

OpenStudy (ray1998):

And if you're not busy, could you help me with some more?

OpenStudy (ray1998):

I could start a new question so that you could get credit for each one :)

rishavraj (rishavraj):

here u go :))) http://prntscr.com/ajoywq

OpenStudy (ray1998):

Oh! Yes :) I got it ^.^ but thank you for screenshotting it to show me.

rishavraj (rishavraj):

u welcome :))